OpenStudy (anonymous):

if g(x) = x2 − 5, find g(7)

OpenStudy (anonymous):

plug in 7 for x

OpenStudy (cathyangs):

^that person is correct! :D

OpenStudy (anonymous):

7^2 - 5 = 49 - 5 = 44

OpenStudy (anonymous):

why you supposed that x2 means x^2 ? ;)

OpenStudy (anonymous):

because try typing x2 in wolfram alpha. it automatically says it is x^2

OpenStudy (cathyangs):

isn't the standardization that you have the coefficient first, for all variables?

OpenStudy (anonymous):

thank you!